OpenStudy (anonymous):

help me please! it saus evaluate the expression -(pq)^2 divided by (-8) for p=2 amd q=4

OpenStudy (anonymous):

=(-(pq)^2)/(-8) =(-(2x4)^2)/(-8) =(-(8)^2)/(-8) =(-64)/(-8) =8

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

ur welcome..., follow my twitter please @gerryphysicist ^^V